# Temporal Overlapping Prediction: A Self-supervised Pre-training Method for LiDAR Moving Object Segmentation

### v1

@@ Abstract
Moving object segmentation (MOS) on LiDAR point clouds is crucial for autonomous systems like self-driving vehicles. Previous supervised approaches rely heavily on costly manual annotations, while LiDAR sequences naturally capture temporal motion cues that can be leveraged for self-supervised learning. In this paper, we propose Temporal Overlapping Prediction (TOP), a self-supervised pre-training method that alleviate the labeling burden for MOS. TOP explores the temporal overlapping points that commonly observed by current and adjacent scans, and learns spatiotemporal representations by predicting the occupancy states of temporal overlapping points. Moreover, we utilize current occupancy reconstruction as an auxiliary pre-training objective, which enhances the current structural awareness of the model. We conduct extensive experiments and observe that the conventional metric Intersection-over-Union (IoU) shows strong bias to objects with more scanned points, which might neglect small or distant objects. To compensate for this bias, we introduce an additional metric called $\text{mIoU}_{\text{obj}}$ to evaluate object-level performance. Experiments on nuScenes and SemanticKITTI show that TOP outperforms both supervised training-from-scratch baseline and other self-supervised pre-training baselines by up to 28.77% relative improvement, demonstrating strong transferability across LiDAR setups and generalization to other tasks. Code and pre-trained models will be publicly available upon publication.
@@ 1 Introduction
Moving Object Segmentation (MOS) on LiDAR data aims to segment moving points from the entire LiDAR scan and provides autonomous systems [9, 33, 34, 50, 54] with important clues in localization and dynamic obstacle avoidance, which enables better planning of collision-free trajectories in dynamic scenes. However, per-point manual labeling of moving objects [9, 3, 6] is both costly and labor-intensive. Meanwhile, self-supervised pre-training achieves significant success in Natural Language Processing (NLP) [13, 30, 19, 40, 5], Computer Vision (CV) [51, 17, 18, 14, 41], and 3D Vision [59, 39, 55, 58], enabling the learning of transferable embeddings from unlabeled raw data. This motivates us to pre-train a backbone using raw LiDAR data in a self-supervised manner and then fine-tune it for MOS, which has limited expensive labeled data.
@@
The MOS task heavily relies on temporal information, including the positions of an object at different times and the changing relative positions of an object to its surroundings. In the literature involving temporal information of LiDAR data, an intuitive approach is to utilize LiDAR observation forecasting as a self-supervised pre-training method for MOS. The forecasting task includes point cloud forecasting  [53, 52, 32, 60, 23] and occupancy forecasting [1], which focus on predicting future scans or occupancy maps based on past and current LiDAR scans and poses. However, utilizing forecasting as a pre-training method for MOS presents certain limitations: (1) The forecasting task is more complex and advanced than MOS. Forecasting a static object mainly requires geometric information to reconstruct the unseen parts of the object from a novel sensor view. For moving objects, beyond geometry, it is essential to extract temporal features such as the object’s kinematics and interactions between objects and their surroundings. Hence, relying on forecasting as a pre-training step for MOS could be overly demanding. (2) Existing methods [23, 1] inevitably learn a “flow” when using raw future observations as supervision. This complex “flow” information represents the relationship between past observations’ embeddings and future observations. Current approaches either implicitly learn this flow through simple feature broadcasting [23] or explicitly predict it [1], lacking direct supervision similar to the scene flow task [61, 62, 29]. Thus, the learned “flow” tends to be noisy. Relying on this noisy “flow” for deterministic future observation predictions (as opposed to the generative approach [60]) can lead to network overfitting, thereby reducing its transferability to target downstream tasks.
Figure 1: Illustration of temporal overlapping points in a realistic scenario. The color of points represents the occupancy state, and the index marked on the points indicates the observed time. For the two temporal overlapping points in the past and one in the future, their temporal occupancy states differ from the current occupancy states.
@@
To overcome these two limitations and incorporate temporal information for MOS self-supervised pre-training, we propose Temporal Overlapping Prediction. TOP explores the temporal occupancy states of temporal overlapping points observed by both the current scan and adjacent past or future scans for pre-training. The rationale for this lies in the potential for moving objects to cause the current occupancy states of temporal overlapping points to differ from the states at other observed times, as illustrated in Figure 1. We first pre-process the spatial positions and temporal occupancy states of the overlapping points with beam divergence [31], which is a physical characteristic of LiDAR. This LiDAR nature helps avoid discretization errors caused by manually setting a fixed voxel resolution. Then, we pre-train the encoder by predicting the pre-processed temporal occupancy states, which simplifies the task compared to directly forecasting LiDAR observations. Based on a distinct temporal correlation between LiDAR beams at different times established by the pre-processing, our pre-training factors out either implicit or explicit learning of a noisy “flow”. Additionally, since understanding current states is also important for MOS, we add a reconstruction term for optimization.
We assess the MOS performance on the nuScenes [6] and SemanticKITTI [3] datasets. Previous evaluations of nuScenes MOS in [34, 46] directly convert object attributes to MOS labels. However, these attributes do not precisely reflect the motion state of the object, and not all objects have attributes. Thus, we combine object attributes with object speed calculated based on bounding boxes to obtain refined nuScenes MOS labels. In the experiments, we observe that the conventional metric Intersection-over-Union (IoU) for MOS exhibits a significant bias towards objects that have a larger number of scanned points, potentially overlooking smaller or more distant objects. Thus, we propose an additional object-level evaluation metric $\text{mIoU}_{\text{obj}}$ to take these objects into account.
@@
We conducted few-shot fine-tuning experiments on the nuScenes dataset with different amounts of labeled data. Results show that TOP achieves improvements over the supervised training-from-scratch baseline [33] and outperforms other self-supervised pre-training baselines, such as state-of-the-art (SOTA) LiDAR observation forecasting methods [23, 1] and the occupancy-based 3D pre-training method [4]. The results of our few-shot cross-dataset experiments on the SemanticKITTI dataset further highlight the superior transferability of TOP pre-training compared to various baselines. Additionally, we fine-tune the pre-trained backbone on the nuScenes semantic segmentation task in a few-shot setting, demonstrating the potential generalization capability of TOP across more tasks.
@@ 2 Related Works
Moving Object Segmentation. Existing supervised MOS methods can be grouped into projection-based and voxel-based approaches. The early MOS methods [9, 24, 45, 11] project 3D LiDAR data to range images and embed the range residual images by 2D backbone. These methods suffer from imprecise boundaries when re-projecting 2D outputs to 3D point clouds [34]. Recent methods [46, 33, 34, 50] utilize the sparse convolution backbone to directly embed voxelized point clouds. 4DMOS [33] first employs sparse 4D UNet [12] and makes it the SOTA backbone for MOS. For unsupervised approaches, several scene flow methods [48, 2, 16, 44] decouple the per-point flow to an ego-motion flow for static points and a non-rigid flow for moving points, which provides a byproduct of MOS [33]. However, these approaches using only two adjacent scans, encounter challenges in identifying slowly moving objects [33]. MOTS [25] proposes a label-free method that only works in stationary scenes. 4dNDF [64] encodes the 4D scene into an implicit neural map, which can be used to distinguish the static map and dynamic objects. However, its application in real-time is limited. In addition, M-detector [54] proposes a network-free approach based on the occlusion principle. It requires manual tuning of numerous hyper-parameters when using different LiDARs or deploying in varying scenes. Moreover, to reduce the cost of manual labeling, several methods [10, 28] propose a pipeline that first generates coarse labels automatically and then refines them manually. In contrast to previous works on MOS, we employ self-supervised pre-training to enhance the few-shot MOS performance.
@@
3D Self-supervised Pre-training. Effective self-supervised pre-training methods in NLP [13, 40, 5] and CV [18] have been extended to 3D data. Some methods [59, 15, 7] leverage BERT-like [13] or GPT-like [40, 5] pre-training on object CAD point clouds. Contrastive learning methods [55, 42, 38, 22, 27, 8] employ various transformations or augmentations to construct point/instance/scene-level contrastive loss, learning geometric or semantic information without manual labels. Additionally, Masked Auto-Encoder (MAE)-based methods [39, 63, 20, 57, 47, 36, 26, 56] extend MAE [18] to 3D point clouds. These methods train an encoder by predicting a certain ratio of masked point cloud patches, which transfers well on object detection and semantic segmentation tasks. Subsequent works [21, 65, 58] apply NeRF-like [35] volumetric neural rendering to leverage multi-modal inputs and bridge the gap between 2D and 3D data. In summary, existing 3D self-supervised pre-training works lack temporal information and thus are unsuitable for MOS.
@@
LiDAR Observation Forecasting. Existing works that involve temporal information of LiDAR data mainly focus on LiDAR observation forecasting, including point cloud and occupancy forecasting. Previous works on point cloud forecasting [53, 52, 32] directly predict the raw future LiDAR scans based on past ones. A recent study, 4Docc [23], claims the potential of forecasting as a scalable self-supervised task. Instead of directly predicting raw point clouds, 4Docc renders the future point clouds on a predicted occupancy map using known sensor intrinsics and extrinsics, enabling the network to focus on learning the world rather than sensor characteristics. Another generative approach, Copilot4D [60] introduces the concept of the world model. It tokenizes LiDAR scans and predicts future scans through discrete diffusion. However, the learned world knowledge is difficult to apply to downstream tasks due to the nature of heavy-decoder. ALSO [4] is a single-scan method that reconstructs the occupancy of a local surface and inspires UnO [1]. UnO is the most recent occupancy forecasting method that predicts the binary occupancy class of points sampled along future LiDAR rays. In this paper, we pre-train the backbone using the proposed TOP method instead of future observation forecasting.
@@ 3 Methodology
Figure 2: The overall pipeline. Pre-processing: (a) Coplanarity condition of two beams. (b) If the spatial angle between two beams $\alpha_{i,j}$ exceeds the beam divergence angle $\theta_{\text{dvg}}$, we sample the beams’ intersection point as the temporal overlapping point. The red and blue areas indicate beam divergence. (c) When $\alpha_{i,j}$ is less than $\theta_{\text{dvg}}$, the intersection segment is calculated to sample temporal overlapping points within it. Pre-training: (d) We use a sparse 4D UNet to embed the input scans. For the temporal overlapping prediction objective $\mathcal{L}_{\text{overlap}}$, the occupancy state of a temporal overlapping point is predicted based on its positional encoding and the embedding of the current beam hit-point. Similarly, the occupancy state of the currently sampled point is also predicted for the current reconstruction objective $\mathcal{L}_{\text{recon}}$.
In this section, we introduce the self-supervised pre-training method TOP. Our method aims to pre-train an encoder for the downstream MOS task. The pre-training goal is to predict the occupancy states of temporal overlapping points and to reconstruct the current occupancy. We first introduce notations and problem formulation in Sec. 3.1. Then, we discuss the pre-processing of the temporal overlapping points in Sec. 3.2. Finally, we describe the pre-training objectives in Sec. 3.3.
@@ 3.1 Problem Formulation
Notations. To begin with, we define a LiDAR scan as $\mathbf{P}\in\mathbb{R}^{N\times 3}$, where $N$ is the number of points. We use subscript $i$ to indicate the $i^{th}$ point in $\mathbf{P}$, where each point $\mathbf{p}_{i}\in\mathbb{R}^{3}$ ($i=1,2,...,N$) contains $xyz$ location. $\mathbf{p}_{i}$ can be further described as a laser beam emitted from the sensor origin $\mathbf{a}\in\mathbb{R}^{3}$ with its normalized beam direction $\mathbf{d}_{i}\in\mathbb{R}^{3}$ and range $r_{i}$, leading to $\mathbf{p}_{i}=\mathbf{a}+r_{i}\mathbf{d}_{i}$. Although the laser beam is commonly modeled as an ideal ray, the real LiDAR beam exhibits a diverging radius as it travels, and this is referred to as beam divergence [31]. The beam divergence angle, denoted as $\theta_{\text{dvg}}$, is a constant for each type of LiDAR. When any part of the diverging beam hits an object, the LiDAR calculates the distance by time of flight and considers the beam center point as the hit-point, as shown in Fig. 3. To incorporate the temporal LiDAR sequence, we indicate scans at different timestamps with superscripts. For example, the scan at the current timestamp $t_{0}$ is $\mathbf{P}^{t_{0}}\in\mathbb{R}^{N\times 3}$. For adjacent past/future timestamps $\mathcal{T}=\{t_{-n}\ldots t_{-1},t_{1}\ldots t_{n}\}$, we have $\mathcal{P}_{\text{adj}}=\{\mathbf{P}^{t}\in\mathbb{R}^{N_{t}\times 3}\}_{t\in\mathcal{T}}$. The $j^{th}$ beam at time $t$, denoted as $\mathbf{d}_{j}^{t}$, can also be described by $\mathbf{p}_{j}^{t}=\mathbf{a}^{t}+r_{j}^{t}\mathbf{d}_{j}^{t}$.
Point Clouds Encoding. All scans in $\mathcal{P}_{\text{adj}}$ are transformed to the local coordinate system of $\mathbf{P}^{t_{0}}$ with the poses provided in the datasets. We take $n$ current and past scans as network inputs $\mathcal{P}_{\text{in}}=\{\mathbf{P}^{t_{-n+1}}\ldots\mathbf{P}^{t_{-1}},\mathbf{P}^{t_{0}}\}$. Each point in $\mathcal{P}_{\text{in}}$ is concatenated with its timestamp as a 4D point before being fed into the network. We apply a sparse 4D UNet encoder $g(\cdot)$ used in [12, 33, 34, 50] to embed the spatiotemporal information of the 4D inputs:
@@
where $\mathbf{F}\in\mathbb{R}^{N\times d}$ is the fully sparse features of the current scan $\mathbf{P}^{t_{0}}$ with a feature dimension of $d$.
Pre-training Loss. We incorporate temporal overlapping points and their temporal occupancy states to pre-train the encoder $g(\cdot)$. As described in Fig. 1, temporal overlapping points are located on current LiDAR beams in $\mathbf{P}^{t_{0}}$ and are also observed by adjacent past/future beams in $\mathcal{P}_{\text{adj}}$. Firstly, we denote the temporal overlapping points as $\mathbf{O}\in\mathbb{R}^{M\times 4}$, where $M$ is the number of points and each point contains $xyz$ position and the temporally observed timestamp $t$. Then, we use the LiDAR occupancy measurement model (introduced in Sec. 3.2) to describe the occupancy state labels of $\mathbf{O}$, denoted as $\mathbf{S}\in\mathbb{R}^{M\times 3}$. Each label is a three-dimensional one-hot vector indicating the occupancy states “free”, “occupied”, and “unknown”. The overall pre-training goal consists of temporal occupancy prediction and current occupancy reconstruction, which is described as:
@@
Figure 3: LiDAR beam divergence model: The bottom-left blue cone illustrates a real diverging beam, where the beam radius expands as the traversed distance increases. When any point within the beam spot hits an obstacle (indicated by the real hit-point colored in red), the center point of the beam spot (colored in black) is considered as the measured point. LiDAR occupancy measurement model: The occupancy state is confidently classified as free before reaching the beam hit-point. Then, the confidence level exponentially decays from the beam hit-point onward. A small area of high confidence after the beam hit-point is categorized as the occupied class, while the remaining areas are deemed unknown.

### v2

@@ Abstract
Moving object segmentation (MOS) on LiDAR point clouds is crucial for autonomous systems such as self-driving vehicles. While previous supervised approaches rely on costly manual annotations, LiDAR sequences naturally capture temporal motion cues that can be leveraged for self-supervised learning. In this paper, we propose Temporal Overlapping Prediction (TOP), a self-supervised pre-training method designed to alleviate this annotation burden. TOP learns powerful spatiotemporal representations by predicting the occupancy states of temporal overlapping points that are commonly observed in current and adjacent scans. To further ground these representations in the current scene’s geometry, we introduce an auxiliary pre-training objective of reconstructing the occupancy of the current scan. Extensive experiments on the nuScenes and SemanticKITTI datasets validate our method’s effectiveness. TOP consistently outperforms existing supervised and self-supervised pre-training baselines across both point-level Intersection-over-Union (IoU) and object-level Recall metrics. Notably, it achieves a relative improvement of up to 28.77% over a training-from-scratch baseline and demonstrates strong transferability across LiDAR setups. Our code is publicly available at https://github.com/ZiliangMiao/TOP.
@@ 1 Introduction
Figure 1: Temporal occupancy inconsistency caused by the relative motion between the sensor and the person. The red dotted lines track fixed points (temporal overlapping points) in space across time, illustrating how their occupancy states change. The point colors denote the occupancy state, while the index indicates the observation time.
@@
Moving Object Segmentation (MOS) on LiDAR data aims to segment moving points from the entire LiDAR scan, providing autonomous systems [9, 33, 34, 50, 55] with important clues for localization and dynamic obstacle avoidance, thereby enabling better planning of collision-free trajectories in dynamic scenes. However, per-point manual labeling of moving objects [9, 3, 6] is both costly and labor-intensive. Meanwhile, self-supervised pre-training has achieved significant success in natural language processing [13, 30, 19, 40, 5], computer vision [51, 17, 18, 14, 41], and 3D vision [60, 39, 56, 59], enabling the learning of transferable embeddings from unlabeled raw data. This motivates us to pre-train a backbone using raw LiDAR data in a self-supervised manner and then fine-tune it for MOS, where labeled data is limited and expensive.
@@
The MOS task heavily relies on temporal information to discern the motion of objects relative to their surroundings. In the literature involving temporal information of LiDAR data, an intuitive approach is to utilize LiDAR observation forecasting as a self-supervised pre-training method for MOS. The forecasting task includes point cloud forecasting  [54, 53, 32, 61, 23] and occupancy forecasting [1], which focus on predicting future scans or occupancy maps based on past and current LiDAR scans and poses. However, utilizing forecasting as a pre-training method for MOS presents certain limitations: (1) The forecasting task is more complex and advanced than MOS. Forecasting a static object mainly requires geometric information to reconstruct the unseen parts of the object from a novel sensor view. For moving objects, beyond geometry, it is essential to extract temporal features such as the object’s kinematics and interactions between objects and their surroundings. Hence, relying on forecasting as a pre-training step for MOS could be unnecessarily complex. (2) Existing methods [23, 1] inevitably learn a “flow” when using raw future observations as supervision. This complex “flow” information represents the relationship between the embeddings of past observations and future observations. Current approaches either implicitly learn this flow through simple feature broadcasting [23] or explicitly predict it [1], lacking direct supervision similar to the scene flow task [62, 63, 29]. Thus, the learned “flow” tends to be noisy. Relying on this noisy “flow” for deterministic future observation predictions (as opposed to the generative approach [61]) can lead to network overfitting, thereby reducing its transferability to the downstream task.
@@
To overcome these two limitations and incorporate temporal information for MOS self-supervised pre-training, we propose TOP, namely Temporal Overlapping Prediction. TOP explores the occupancy states of temporal overlapping points observed by both the current scan and adjacent (past or future) scans for pre-training. The rationale is that the motion of the sensor and objects creates temporal occupancy inconsistency, where a temporal overlapping point may exhibit different occupancy states when observed at different times, as shown in Fig 1. We first pre-process the spatial positions and occupancy states of the temporal overlapping points by considering beam divergence [31], which is a physical characteristic of LiDAR. Leveraging this physical characteristic of LiDAR helps avoid discretization errors caused by manually setting a fixed voxel resolution. Then, we pre-train the encoder by predicting the pre-processed occupancy states, which simplifies the task compared to directly forecasting LiDAR observations. Based on a distinct temporal correlation between LiDAR beams at different times established by the pre-processing, our pre-training avoids the need for either implicitly or explicitly learning a noisy “flow”. Additionally, since understanding the current scene geometry is also important for MOS, we include a reconstruction term in our pre-training objective.
@@
We evaluate TOP via few-shot fine-tuning on nuScenes [6] with varying amounts of labeled data. The results demonstrate that TOP outperforms both the supervised training-from-scratch baseline [33] and other self-supervised pre-training baselines, including state-of-the-art (SOTA) LiDAR forecasting methods [23, 1] and the occupancy-based 3D method [4]. Furthermore, cross-dataset experiments on SemanticKITTI [3] highlight TOP’s superior transferability, while its few-shot performance on the nuScenes semantic segmentation task demonstrates its potential to generalize to other downstream tasks.
@@ 2 Related Works
Moving Object Segmentation. Existing supervised MOS methods can be grouped into projection-based and voxel-based approaches. The early MOS methods [9, 24, 45, 11] project 3D LiDAR data onto range images and embed the range residual images using a 2D backbone. These methods suffer from imprecise boundaries when re-projecting 2D outputs to 3D point clouds [34]. Recent methods [46, 33, 34, 50] utilize a sparse convolution backbone to directly embed voxelized point clouds. 4DMOS [33] first employed sparse 4D UNet [12] and established it as the SOTA backbone for MOS. For unsupervised approaches, several scene flow methods [48, 2, 16, 44] decouple the per-point flow into an ego-motion flow for static points and a non-rigid flow for moving points, which yields MOS as a byproduct. However, these two-frame approaches encounter challenges in identifying slowly moving objects [33]. MOTS [25] proposes a label-free method that is limited to stationary scenes. 4dNDF [65] encodes the 4D scene into an implicit neural map, which can be used to distinguish between the static map and dynamic objects. However, its computational cost limits real-time applicability. In addition, M-detector [55] proposes a network-free approach based on the occlusion principle. It requires manual tuning of numerous hyperparameters when using different LiDARs or deploying in different scenes. To reduce manual labeling costs, several methods [10, 28] propose automatic labeling pipelines, though some of them still incorporate a manual refinement step. In contrast to previous MOS works, we employ self-supervised pre-training to enhance the few-shot MOS performance.
@@
3D Self-supervised Pre-training. Effective self-supervised pre-training methods in natural language processing [13, 40, 5] and computer vision [18] have been extended to 3D data. Some methods [60, 15, 7] leverage BERT-like [13] or GPT-like [40, 5] pre-training on object CAD point clouds. Contrastive learning methods [56, 42, 22, 27, 8] employ various transformations or augmentations for contrastive losses at point, instance, and scene levels, learning geometric or semantic information without manual labels. Inspired by MAE [18], numerous methods [39, 64, 20, 58, 47, 36, 26, 57] pre-train an encoder by reconstructing masked point cloud patches, which yields effective 3D representations for object detection and semantic segmentation. Subsequent works [21, 66, 59] apply NeRF-like [35] volumetric neural rendering to integrate 2D and 3D multi-modal data. To the best of our knowledge, while a few methods [38, 52] incorporate temporal information, TOP is the first self-supervised pre-training method specifically designed for dynamic tasks like MOS.
@@
LiDAR Observation Forecasting. Existing works that involve temporal information of LiDAR data mainly focus on LiDAR observation forecasting, including point cloud and occupancy forecasting. Previous works on point cloud forecasting [54, 53, 32] directly predict the future LiDAR scans based on past ones. A recent study, 4DOcc [23], highlights the potential of forecasting as a scalable self-supervised task. Instead of directly predicting raw point clouds, 4DOcc renders the future point clouds on a predicted occupancy map using known sensor intrinsics and extrinsics, enabling the network to focus on learning the world rather than sensor characteristics. Another generative approach, Copilot4D [61], introduces the concept of the world model. It tokenizes LiDAR scans and predicts future scans through discrete diffusion. However, the learned world knowledge is difficult to apply to downstream tasks due to its heavy-decoder architecture. ALSO [4] is a single-scan method that reconstructs the occupancy of a local surface. This work inspired UnO [1], which is the most recent occupancy forecasting method that predicts the binary occupancy class of points sampled along future LiDAR rays. In contrast, our work diverges from these forecasting-based approaches by pre-training the backbone with our proposed TOP method.
@@ 3 Methodology
Figure 2: LiDAR beam divergence: The diverging beam is illustrated as the blue cone. The reported point (black) lies on the beam’s centerline, while the actual beam hit-point (red) can be anywhere within the beam’s footprint. LiDAR occupancy measurement: We model the space from the sensor to the reported LiDAR point as free with maximum confidence. Beyond this point, confidence decays exponentially. The region immediately following the reported point is considered occupied if its confidence is above a predefined threshold. The subsequent region where the confidence falls below the threshold is considered unknown.
Figure 3: Overall pipeline. Pre-processing: (a) shows the coplanarity condition of two beams. (b) If the spatial angle between two beams $\alpha_{i,j}$ exceeds the beam divergence angle $\theta_{\text{dvg}}$, we sample the beams’ intersection point as the temporal overlapping point. The red and blue areas indicate beam divergence. (c) When $\alpha_{i,j}$ is less than $\theta_{\text{dvg}}$, we calculate the intersection segment and sample temporal overlapping points from it. Pre-training: (d) The input sequence is encoded by a sparse 4D UNet. For both pre-training objectives, a shallow MLP decoder predicts occupancy states based on the point’s positional encoding and the feature of its corresponding LiDAR point.
In this section, we introduce TOP, a self-supervised method to pre-train an encoder for the downstream MOS task. The pre-training goal involves two objectives: predicting the occupancy of temporal overlapping points and reconstructing the current scene’s occupancy. We first introduce the notation and problem formulation in Sec. 3.1. Then, we discuss the pre-processing of the temporal overlapping points in Sec. 3.2. Finally, we describe the pre-training objectives in Sec. 3.3.
@@ 3.1 Problem Formulation
Notations. To begin with, we define a LiDAR scan as $\mathbf{P}\in\mathbb{R}^{N\times 3}$, where $N$ is the number of points. We use subscript $i$ to indicate the $i^{th}$ point in $\mathbf{P}$, where each point $\mathbf{p}_{i}\in\mathbb{R}^{3}$ ($i=1,2,\dots,N$) contains xyzxyz coordinates. $\mathbf{p}_{i}$ can be further described by a laser beam emitted from the sensor origin $\mathbf{a}\in\mathbb{R}^{3}$ with its normalized beam direction $\mathbf{d}_{i}\in\mathbb{R}^{3}$ and range $r_{i}$, leading to $\mathbf{p}_{i}=\mathbf{a}+r_{i}\mathbf{d}_{i}$. Although the laser beam is commonly modeled as an ideal ray, an actual LiDAR beam exhibits a diverging radius as it travels, a phenomenon known as beam divergence [31]. The beam divergence angle, denoted as $\theta_{\text{dvg}}$, is a constant for each type of LiDAR.
When a diverging LiDAR beam hits an obstacle and reflects, the sensor reports a point on the beam’s centerline corresponding to the measured time-of-flight range. In contrast, the actual beam hit-point (reflection point) can be located anywhere within the beam’s footprint, as shown in Fig. 2. To incorporate the temporal LiDAR sequence, we indicate scans at different timestamps with superscripts. For example, the scan at the current timestamp $t_{0}$ is $\mathbf{P}^{t_{0}}\in\mathbb{R}^{N_{t_{0}}\times 3}$. For adjacent past/future timestamps $\mathcal{T}=\{t_{-n},\dots,t_{-1},t_{1},\dots,t_{n}\}$, we have $\mathcal{P}_{\text{adj}}=\{\mathbf{P}^{t}\in\mathbb{R}^{N_{t}\times 3}\}_{t\in\mathcal{T}}$. The $j^{th}$ beam at time $t$, $\mathbf{d}_{j}^{t}$, results in the point $\mathbf{p}_{j}^{t}$, which is described by $\mathbf{p}_{j}^{t}=\mathbf{a}^{t}+r_{j}^{t}\mathbf{d}_{j}^{t}$.
Point Cloud Encoding. All scans in $\mathcal{P}_{\text{adj}}$ are transformed to the local coordinate system of $\mathbf{P}^{t_{0}}$ using the poses provided in the datasets. We take $n$ current and past scans as network inputs $\mathcal{P}_{\text{in}}=\{\mathbf{P}^{t_{-n+1}},\dots,\mathbf{P}^{t_{-1}},\mathbf{P}^{t_{0}}\}$. Each point in $\mathcal{P}_{\text{in}}$ is concatenated with its timestamp before being fed into the network. We apply a sparse 4D UNet encoder $g(\cdot)$ used in [12, 33, 34, 50] to embed the spatiotemporal information of the 4D inputs:
@@
where $\mathbf{F}\in\mathbb{R}^{N_{t_{0}}\times d}$ is a sparse feature for points in the current scan $\mathbf{P}^{t_{0}}$ with a feature dimension of $d$.
Pre-training Loss. We incorporate temporal overlapping points and their temporal occupancy states to pre-train the encoder $g(\cdot)$. As described in Fig. 1, temporal overlapping points are locations along the LiDAR beams from the current scan $\mathbf{P}^{t_{0}}$ that are also intersected by beams from adjacent past/future scans $\mathcal{P}_{\text{adj}}$. Firstly, we denote the temporal overlapping points as $\mathbf{O}\in\mathbb{R}^{M\times 4}$, where $M$ is the number of points and each point has xyzxyz coordinates and the timestamp $t$ of the adjacent scan that observed it. Then, we use the LiDAR occupancy measurement (introduced in Sec. 3.2) to determine the occupancy state labels for $\mathbf{O}$, denoted as $\mathbf{S}\in\mathbb{R}^{M\times 3}$. The overall pre-training loss $\mathcal{L}$, further described in Sec. 3.3, consists of temporal overlapping prediction $\mathcal{L}_{\text{overlap}}$ and current occupancy reconstruction $\mathcal{L}_{\text{recon}}$: